OpenStudy (anonymous):

Solve the following system of equations. x + 3y – z = 2 x – 2y + 3z = 7 x + 2y – 5z = –21 A. (–2, 3, 5) B. (2, 3, –5) C. (2, –3, 5) D. (2, 3, 5)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

sav u must do it for ur self

OpenStudy (anonymous):

i honestly don't know how.

OpenStudy (anonymous):

i know you can

OpenStudy (anonymous):

you can try solving by substitution

OpenStudy (anonymous):

or try this, eq 1 mult by 3 to cancel out z, by adding eq 2 and eq1 x + 3y – z = 2 eq (1) mult by 3 to get 3x + 9y -3z = 6 x – 2y + 3z = 7 eq(2) +(x – 2y + 3z = 7) x + 2y – 5z = –21 eq (3) -------------------- 4x+7y+0=13 eq 4 mult eq1 by 5 then subtract to eq3 5x+15y-5z= 10 -(x + 2y – 5z = –21) eq (3) -------------- 4x+13y+0 =31 eq5 subtract this to eq 4 -(4x+7y+0=13) eq 4 --------------- 0 +6y =18 y=18/6=3 ans now sub this to eq(4 or 5) 4x+7y+0=13 eq 4 4x+7(3)+0=13 eq 4 x=-2 ans then sub x=-2 and y=3 to eq1 x + 3y – z = 2 -2 + 3(3) – z = 2 z=5 ans therefore (x,y,z)=(-2,3,5) answer